OpenStudy (lucaz):

x=[0, 2pi] find the sum of the roots of the equation:

OpenStudy (lucaz):

\[4\left( \cos x \right)^{3}-3\cos x=0\]

OpenStudy (campbell_st):

well factor out cos(x) and you have \[\cos(x)(4\cos^2(x) - 3) = 0\] so you now need to factor the quadratic and you get \[\cos(x)(2\cos(x) - \sqrt{3})(2\cos(x) + \sqrt{3}) = 0\] I'll leave you to find the values of x in the given domain that make the equation true

OpenStudy (lucaz):

okay

OpenStudy (campbell_st):

do you know what to do next...?

OpenStudy (lucaz):

I think i have to find 2cos(x)=sqrt(3) ?

OpenStudy (anonymous):

4(cosx)^3=3(cosx) 4(cosx)^2=3 (Cosx)^2=(3/4) Cosx=\[\pm \sqrt{3/4}\] Cos x = \[\frac{ \sqrt{3} }{2 }\]

OpenStudy (campbell_st):

lol... there you go.... why worry about learning where people just post solutions... but there is a missing solution...

OpenStudy (lucaz):

I found this sqrt(3)/2 but wasn't sure what that meant

OpenStudy (campbell_st):

well thats correct as posted by @aj.est1979 there is a solution where \[\cos(x) = \pm \frac{\sqrt{3}}{2}\] there are exact value angles... so easy to find.... go back to my factorization to find the other value cos(x) can take

OpenStudy (campbell_st):

but remember... you need angles...

OpenStudy (lucaz):

cos(x)=0, x=pi/2?

OpenStudy (lucaz):

really helpful

OpenStudy (campbell_st):

that correct @lucaz it has cos(x) = 0 and in the given domain, cos(x) = 0 for 2 angles... one is \[\frac{\pi}{2}\] and there is another...

OpenStudy (lucaz):

I think it's 3pi/2

OpenStudy (campbell_st):

thats correct... so now you have all the solutions....

OpenStudy (lucaz):

hey, this kind of problem the domain is always ginen with pi, or it could be in degrees?

OpenStudy (campbell_st):

thats correct... your answers are expressed as radians.... if the domain was degrees just change them to radians.. pi/2 = 90 3pi/2 = 270

OpenStudy (lucaz):

right, thank you